What is the slope given 2x-y=3
Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

What is the slope given 2x-y=3
--------
Add y to both sides to get:
y+3 = 2x
---
Solve for "y":
y = 2x-3
----
slope = 2
--------------
Cheers,
Stan H.
==============

Answer by Cromlix(4381)

First of all sort 2x - y = 3
into the form: y = mx + c
-y = -2x + 3
Multiply both sides by -1
y = 2x - 3
Slope = 2
Intersection with y axis = 0,-3
Hope this helps.
:-)